# RWKV-UNet: Improving UNet with Long-Range Cooperation for Effective Medical Image Segmentation

### v1

@@ 3.4 Cross-Channel Mix for Multi-scale Fusion
Define output feature map of Stage \@slowromancapi@, Stage \@slowromancapii@ and Stage \@slowromancapiii@ of the designed encoder are $F_{1},F_{2}$, and $F_{3}$, which has different Spatial dimensions and channel counts: $F_{1}\in\mathbb{R}^{C_{1}\times H_{1}\times W_{1}},\quad F_{2}\in\mathbb{R}^{C_{2}\times H_{2}\times W_{2}},\quad F_{3}\in\mathbb{R}^{C_{3}\times H_{3}\times W_{3}}$. ($H_{1}$ >$H_{2}$>$H_{3}$, $W_{1}$ >$W_{2}$ >$W_{3}$ and $C_{3}$ >$C_{2}$ >$C_{1}$).
@@
where $\tilde{F}_{1}\in\mathbb{R}^{C_{1}\times H_{1}\times W_{1}}$, $\tilde{F}_{2}\in\mathbb{R}^{C_{1}\times H_{1}\times W_{1}}$, and $F_{3}\in\mathbb{R}^{C_{1}\times H_{1}\times W_{1}}$.
@@
where $F_{1}^{\prime}\in\mathbb{R}^{C_{1}\times H_{1}\times W_{1}}$ , $F_{2}^{\prime}\in\mathbb{R}^{C_{2}\times H_{2}\times W_{2}}$ and $F_{3}^{\prime}\in\mathbb{R}^{C_{3}\times H_{3}\times W_{3}}$. $F_{1}^{\prime}$,$F_{2}^{\prime}$ and $F_{3}^{\prime}$ will be concatenated with the output of the decoder Stage \@slowromancapiii@, Stage \@slowromancapii@ and Stage \@slowromancapi@.
4 Experiments
Figure 2: Average DSC vs. #Params for different methods on Synapse multi-organ segmentation dataset.
Figure 3: Average DSC vs. #FLOPs for different methods on Synapse multi-organ segmentation dataset.
[…]
@@ 4 Experiments
RWKV-UNet-T (ours) 17.83 81.58 89.28 64.18 86.03 83.76 95.09 63.50 90.36 80.44
RWKV-UNet-S (ours) 21.16 82.20 89.36 67.60 84.11 79.26 95.60 68.86 91.58 81.21
RWKV-UNet (ours) 15.70 84.02 89.53 68.94 87.63 84.07 95.57 69.38 90.95 86.09
Table 4: Comparison results on ACDC dataset. The evaluation metric is DSC in (%).
@@
RWKV-UNet-T (ours) 90.92 89.73 87.18 95.85
RWKV-UNet-S (ours) 91.64 89.77 88.48 96.65
RWKV-UNet (ours) 92.17 90.86 88.72 96.92
@@
Table 5: Comparison results on BUSI, CVC-ClinicDB, CVC-ColonDB, Kvasir-SEG, ISIC 2017 and GLAS datasets. The evaluation metric are DSC in (%). Results with error bars are averaged over three runs with three different splits.
@@
Methods BUSI CVC-ClinicDB CVC-ColonDB Kvasir-SEG ISIC 2017 GLAS Params. FLOPs
U-Net [46] 74.90$\pm$1.15 89.22$\pm$4.07 83.43$\pm$5.24 84.94$\pm$2.60 82.94 88.56$\pm$0.63 07.8M 12.15G
Att-UNet [40] 78.14$\pm$1.25 90.86$\pm$4.09 84.75$\pm$6.55 86.21$\pm$2.15 83.96 89.14$\pm$0.54 08.7M 16.79G
UNet++ [61] 75.74$\pm$1.70 90.07$\pm$5.09 84.75$\pm$3.99 85.86$\pm$1.52 82.27 88.93$\pm$0.67 09.2M 34.77G
TransUNet [9] 79.35$\pm$2.45 93.91$\pm$2.28 90.72$\pm$3.92 91.34$\pm$0.01 86.09 91.58$\pm$0.15 00.1G 38.54G
FCBFormer [48] 80.66$\pm$0.28 91.15$\pm$3.50 85.23$\pm$5.21 91.09$\pm$0.96 84.56 91.13$\pm$0.25 53.0M 41.60G
PVT-CASCADE  [44] 78.30$\pm$3.17 93.45$\pm$1.67 90.97$\pm$2.11 91.40$\pm$1.29 84.77 91.85$\pm$0.39 35.3M 08.40G
Rolling-UNet-M  [31] 77.49$\pm$2.42 91.91$\pm$2.56 85.20$\pm$5.85 87.41$\pm$1.78 84.42 88.12$\pm$0.66 07.1M 08.31G
PVT-EMCAD-B2  [45] 79.22$\pm$1.07 93.39$\pm$1.64 91.27$\pm$1.84 90.37$\pm$1.10 84.05 91.77$\pm$0.09 26.8M 05.85G
U-KAN  [30] 76.95$\pm$1.67 89.88$\pm$3.21 84.66$\pm$5.33 85.78$\pm$1.68 83.46 87.42$\pm$0.28 25.4M 10.59G
RWKV-UNet-T (ours) 79.47$\pm$1.35 95.05$\pm$0.67 89.56$\pm$5.26 91.26$\pm$2.13 85.59 91.96$\pm$0.31 03.3M 03.78G
RWKV-UNet-S (ours) 80.92$\pm$1.74 95.58$\pm$0.69 91.30$\pm$3.21 91.78$\pm$1.24 86.38 92.11$\pm$0.16 09.9M 07.75G
RWKV-UNet (ours) 81.92$\pm$0.82 95.26$\pm$0.70 92.27$\pm$2.48 91.26$\pm$1.27 85.32 92.35$\pm$0.07 17.4M 14.64G
4.1 Experimental Setup
Datasets and Metrics.
@@ 4.1 Experimental Setup
Experiments on conducted on Synapse [27] for multi-organ Segmentation in CT images, ACDC [5] for cardiac segmentation in MRI images, BUSI [1] for breast tumor segmentation in ultrasound images, CVC-ClinicDB [4], CVC-ColonDB [3], Kvasir-SEG [25] for poly segmentation, ISIC 2017 [13] for skin lesion segmentation and GLAS  [49] for gland segmentation. The average Dice Similarity Coefficient (DSC) and average 95% Hausdorff Distance (HD95) are used as evaluation metrics. Specific details for datasets can be seen in Appendix.
Comparison Methods.
We use U-Net  [46], R50 U-Net  [20, 46], R50 Att-Unet  [20, 40], ViT  [15] + CUP  [9], R50-ViT  [20, 15] + CUP  [9], TransUNet  [9], SwinUNet  [6], TransClaw U-Net  [7], MT-UNet  [53], UCTransNet  [52], MISSFormer  [23], CASTformer  [57], TransDeepLab  [2], CTC–Net  [58], MS-UNet  [8], LeVit-Unet-384  [55], HiFormer-L  [21], PVT-CASCADE  [44], TransCASCADE  [44], AiA-UNet  [43], MetaSeg-B  [26], VM-UNet  [47],PVT-EMCAD-B2  [45], SelfReg-UNet  [63, 46], SelfReg-SwinUnet  [63, 6], FCBFormer [48], Rolling-UNet-M  [31], U-KAN  [30]as baselines.
Implementation Details.
All experiments are all performed on PG500-216(V-100) with 32 GB memory. The resolution of the input images is resized to 224$\times$224 for Synapse and ACDC, and 256$\times$256 for binary segmentation tasks. The total training epochs for Synapse are 30, for ACDC are 150, and for binary segmentation are 300. The batch size is 24 for Synapse and ACDC, and 8 for binary segmentation tasks. The initial learning rate is 1e-3 for Synapse, 5e-4 for ACDC, and 1e-4 for binary segmentation experiments. The minimum learning rates for Synapse and ACDC are 0 and for binary segmentation are 1e-5. The AdamW  [36] optimizer and CosineAnnealingLR  [37] scheduler are used. Test-time augmentation is used for Synapse and ACDC. The baseline results on binary segmentation tasks are run by ourselves. The loss function used is a mixed loss combining cross entropy (CE) loss and dice loss  [39]:
@@
where $\alpha$ and $\beta$ are 0.5 and 0.5 for Synapse, 0.3 and 0.7 for ACDC, 0.5 and 1 for binary segmentation. Experiments for PVT-CASCADE  [44] and PVT-EMCAD-B2  [45] use deep supervision strategy  [29].
@@
4.2 Comparison with State-of-the-arts
Multi-organ Segmentation.
Table 3 shows that RWKV-UNet excels in abdomen organ segmentation on the Synapse multi-organ dataset, achieving the highest average DICE score of 84.02% and surpassing all SOTA CNN- and transformer-based methods. It also achieves 15.70 in HD95, only slightly lower than PVT-EMCAD-B2, demonstrating a strong ability to localize the organ boundary. Its performance on the most challenging task, pancreas segmentation, is the best, indicating its effectiveness in handling sample imbalance issues. Smaller models, RWKV-UNet-T and RWKV-UNet-S, also achieve remarkable results.
Cardiac Segmentation.
Table 4 shows that our model outperforms all SOTA models in the ACDC MRI cardiac segmentation task, achieving the best results in both RV and LV categories. The smaller models, RWKV-UNet-T and RWKV-UNet-S, also perform well.
Binary Segmentation.
Table 5 shows that RWKV-UNet series show remarkable performance in the tasks of breast lesion, poly, skin lesion and grand segmentation, with much smaller parameters and computation load than TransUNet and FCBFormer. Meanwhile, it is observed that in some datasets the smaller model RWKV-UNet-S perform better.
4.3 Ablation and Analysis
Effect of Pre-training for Encoder.
Whether to use pre-trained weights of RWKV-UNet encoder results in Table 6. Experimental results indicate that pre-training is essential because it significantly improves feature extraction by enabling the encoder to capture intricate patterns and hierarchical information. Notably, without pre-training, HD95 is higher. But DSC is a more important metric, providing a reliable assessment of segmentation quality directly, while HD95 can be overly sensitive to outliers.
Table 6: Comparison results on Synapse dataset of whether to use pre-trained weights on ImageNet-1k or not.
@@ 4.3 Ablation and Analysis
w/o 30 13.94 76.49
w/o 60 14.03 78.68
w/ 30 15.70 84.02
Table 7: Comparison results of different decoder designs for RWKV-UNet on Synapse dataset. $c\rightarrow c\rightarrow d$ denotes $\operatorname{Conv}_{1\times 1}\rightarrow\operatorname{Conv}_{1\times 1}\rightarrow$ DW-Conv and $c\rightarrow d\rightarrow c$ denotes $\operatorname{Conv}_{1\times 1}\rightarrow$ DW-Conv $\rightarrow\operatorname{Conv}_{1\times 1}.$
Decoder Kernel Size Average Flops
@@
Conv + upsample 3 24.95 81.77 11.38G
$c\rightarrow c\rightarrow d$
+ upsample
3 23.44 82.10 11.06G
5 25.53 81.84 11.07G
7 20.78 81.57 11.08G
9 18.07 83.76 11.10G
$c\rightarrow d\rightarrow c$
+ upsample
3 21.68 83.33 11.08G
5 22.26 83.74 11.11G
7 22.90 82.54 11.16G
9 15.70 84.02 11.22G
Decoder Design.
In Tabel 7, we examine various sizes of convolution kernels, finding that the use of large kernels, such as $\times 9$ 9, in the decoder significantly enhances the segmentation performance. This improvement is likely due to larger kernels capturing broader contextual information and preserving more spatial details, which enhances segmentation accuracy in medical images. We also compare different ways to construct the decoder block and the results show that $\operatorname{Conv}_{1\times 1}\rightarrow\operatorname{Conv}_{1\times 1}\rightarrow$ DW-Conv performs better than $\operatorname{Conv}_{1\times 1}\rightarrow$ DW-Conv $\rightarrow\operatorname{Conv}_{1\times 1}$ and the common convolution layer, showing that DW-Conv operating on larger dimensions will help the decoding process.
Channel Mix.
In VRWKV [16], a Channel Mix module is used after the Spatial Mix. In RWKV-UNet’s encoder, IR-RWKV blocks are designed without using Channel Mix as the pointwise convolution in the output layer performs both channel mix and feed-forward operations. We compare the results on the Synapse dataset with and without applying channel mix after spatial mix. All experiments are performed without pre-training, with a total of 60 epochs. The results are shown in Table 8, demonstrating that the channel mix is not essential for maintaining performance.
Table 8: Comparison of results on the Synapse dataset: effect of applying Channel Mix after Spatial Mix in the IR-RWKV blocks
@@
w/ 15.29 78.76 19.49G
w/o (ours) 14.03 78.68 11.22G
Skip Connections.
Comparison results of different skip connection designs for RWKV-UNet on the Synapse dataset are shown in Table 9. Experiments analyze variations in the number of skip connections and the inclusion of the CCM block. We prioritize retaining shallow skip connections because shallow layers tend to experience greater information loss. The results show that, like other UNet’s variants, adequate skip connections are essential for RWKV-UNet’s performance. Furthermore, the results show that the CCM module can significantly improve the model segmentation results. Even if only one or two inputs from the encoder(when only one input, the CCM block degrades to a ”channel mix bridge”), the CCM block can still lead to increments in DSC. Admittedly, the CCM module increases the computational load to some extent due to the aggregation of global information.
Table 9: Comparison results on different skip connections design of RWKV-UNet on Synapse dataset. Experiments are conducted on the number of skips and whether or not to use the CCM block.
@@
0 w/o 23.39 74.19 09.25 G
1 w/o 17.17 79.76 09.41 G
1 w/ 19.56 79.98 09.63 G
2 w/o 16.08 80.90 09.49 G
2 w/ 19.97 81.11 10.24 G
3 w/o 22.72 81.14 09.56 G
3 w/ 15.70 84.02 11.22 G
5 Conclusion
In this study, we introduce RWKV-UNet, a novel architecture that integrates the RWKV architecture with U-Net. By combining the strengths of convolutional networks for local feature extraction and RWKV’s ability to model global context, our model significantly improves medical image segmentation accuracy. The proposed enhancements, such as the IR-RWKV module and the CCM module, contribute to a more precise feature representation and information fusion across different scales. Experimental results on benchmark datasets demonstrate that RWKV-UNet surpasses all SOTA CNN- and transformer-based methods. Its variants (RWKV-UNet-S and RWKV-UNet-T) offer a practical balance between performance and computational efficiency. Our approach has a strong potential to advance medical image analysis, particularly in clinical settings where both accuracy and efficiency are paramount.
Limitations and Future Work. RWKV-UNet is a powerful 2D medical image segmentation model that effectively combines RWKV with convolution, but it is currently not applicable to 3D imaging. In the future, we plan to extend the model to 3D for handling volume data and to explore the potential of RWKV-based foundational models for medical imaging. We also aim to develop ultra-lightweight RWKV-based models tailored for ’bedside healthcare’ applications, preserving segmentation accuracy while enhancing adaptability and speed further.
Appendix
The supplementary material gives more specific details of our experimental settings, and presents more comprehensive analysis and results of our RWKV-UNet to facilitate the comparison of subsequent methods:
  • Sec. A provides introduction to datasets and corresponding splitting settings used in this paper.
  • Sec. B.1 provides comparison results on Synapse dataset to explore the effect of applying DW-Conv after Spatial Mix in IR-RWKV blocks.
  • Sec. B.2 provides comparison results of different methods on Synapse dataset with larger inputs.
Appendix A Introduction to datasets
1) The Synapse multi-organ segmentation dataset  [27]: 30 abdominal CT scans from the MICCAI 2015 Multi-Atlas Abdomen Labeling Challenge are used. The dataset is split into 18 training cases and 12 cases for validation following the setting in  [9].
2) The Automated Cardiac Diagnosis Challenge (ACDC)  [5]: This dataset gathezrs examinations obtained by MRI scanners. Each patient scan is manually annotated with ground truth data for the left ventricle (LV), right ventricle (RV), and myocardium (MYO). The dataset is split into 70 training cases, along with 10 cases for validation and 20 for testing following the setting in  [9].
3) Breast Ultrasound Images Dataset (BUSI) [1]: This dataset contains breast ultrasound images collected from women between 25 and 75 years old. The dataset consists of 780 pairs of images in normal, benign, and malignant, with corresponding masks. Unlike the setting in [50], we include all three classes in the dataset. The dataset is split into the training set, the validation set and the validation set with a ratio of 8:1:1 with three different random states.
4) CVC-ClinicDB  [4]:This dataset contains frames extracted from colonoscopy videos that contain several examples of polyps. The dataset consists of 612 pairs of images and corresponding masks. he dataset is split into the training set, the validation set and the testing set with a ratio of 8:1:1 with three different random states.
5) CVC-ColonDB [3]: This dataset includes 380 images, each with a resolution of 574$\times$500 pixels. These images were selected from a pool of 12,000 images obtained from 15 short colonoscopy videos, with only these 380 images being annotated. The dataset is split into the training set, the validation set and the testing set with a ratio of 8:1:1 with three different random states.
6) Kvasir-SEG  [25]: This dataset comprises 1000 polyp images and their corresponding ground truth masks with resolutions ranging from 332$\times$487 to 1920$\times$1072 pixels. he dataset is split into the training set, the validation set and the testing set with a ratio of 8:1:1 with three different random states.
7) ISIC 2017  [13]: This dataset contains skin lesions images for Skin Lesion Analysis toward Melanoma Detection Challenge at the International Symposium on Biomedical Imaging (ISBI) 2017. There are 2000 pairs of images and corresponding masks in the training set, 150 pairs in the validation set, and 600 pairs in the testing set.
8) GlaS  [49]: This dataset is designed for gland segmentation. It consists of 165 whole-slide images (WSIs), with 85 images designated for training and 80 for testing. The training dataset is split into a new training set and a validation set with a ratio of 9:1 with three different random states.
Appendix B More Ablation and Exploration Analysis
B.1 Depth-wise Convolution in IR-RWKV Blocks
The comparison results to explore the effect of DW-Conv after Spatial Mix in IR-RWKV blocks are shown in Table A1. The inclusion of DW-Conv demonstrates a notable improvement in performance metrics. The slight increase in computational cost is minimal and well-justified given the significant gains in performance. This indicates that in IR-RWKV blocks, DW-Conv plays a critical role. It provides local information that complements the long-range dependencies captured by Spatial Mix, effectively enhancing model accuracy. Without pre-training, relying solely on Spatial Mix without the inclusion of local information from DW-Conv limits the model’s ability to fully capture fine-grained spatial details. This demonstrates the necessity of integrating both components to achieve a balanced representation of local and global features.
Table A1: Comparison of results on the Synapse dataset: effect of applying DW-Conv after Spatial Mix in IR-RWKV blocks
DW-Conv Average FLOPs
@@ B.1 Depth-wise Convolution in IR-RWKV Blocks
w/o 16.55 75.56 11.14G
w/ (ours) 14.03 78.68 11.22G
B.2 Larger Resolution Inputs
The comparison results of different methods with 512$\times$ 512 input in the Synapse dataset are shown in Table A2. The average Dice scores of other methods are from  [10]. Our results indicate that when the input resolution is increased to 512$\times$512, our model demonstrates an improvement in performance. Although the improvement is not significant, it still maintains the best performance overall. Furthermore, compared to TransUNet, the computational overhead associated with the resolution increase is less severe, showcasing the efficiency of our model at higher resolutions.
Figure A1: #FLOPs vs input resolutions. for different methods. The input channel is 1 and the output is 9 (aligned with settings on Synapse dataset).
Table A2: Comparison of different methods with 512$\times$512 resolution input on the Synapse dataset. The evaluation metric are DSC in (%). The average DSC of other methods are from  [10].
@@ B.2 Larger Resolution Inputs
TransUNet  [9] 84.36 168.7G
RWKV-UNet (ours)  [9] 85.52 058.6G
As shown in Figure A1, the RWKV-UNet series demonstrates advantages in handling high-resolution inputs, as shown in the FLOPs comparison. Particularly for larger resolutions, the RWKV-UNet models maintain significantly lower computational costs compared to other high-performance architectures like transUNet and MT-UNet. It demonstrates that linear attention outperforms traditional Transformers in efficiency, enabling faster and more memory-efficient processing of high-resolution inputs and long sequences.

### v2

@@ III-D Cross-Channel Mix for Multi-scale Fusion
Define output feature map of Stage I, Stage II and Stage III of the designed encoder are $F_{1},F_{2}$, and $F_{3}$, which has different Spatial dimensions and channel counts: $F_{1}\in\mathbb{R}^{C_{1}\times H_{1}\times W_{1}},\quad F_{2}\in\mathbb{R}^{C_{2}\times H_{2}\times W_{2}},\quad F_{3}\in\mathbb{R}^{C_{3}\times H_{3}\times W_{3}}$. ($H_{1}$ $H_{2}$$H_{3}$, $W_{1}$ $W_{2}$ $W_{3}$ and $C_{3}$ $C_{2}$ $C_{1}$).
@@
where $\tilde{F}_{i}\in\mathbb{R}^{C_{1}\times H_{1}\times W_{1}},\quad i=1,2,3$
@@
where $F_{1}^{\prime}\in\mathbb{R}^{C_{1}\times H_{1}\times W_{1}}$ , $F_{2}^{\prime}\in\mathbb{R}^{C_{2}\times H_{2}\times W_{2}}$ and $F_{3}^{\prime}\in\mathbb{R}^{C_{3}\times H_{3}\times W_{3}}$. $F_{1}^{\prime}$$F_{2}^{\prime}$ and $F_{3}^{\prime}$ will be concatenated with the output of the decoder Stage III, Stage II and I.
IV Experiments
TABLE III: Comparison results on Synapse dataset. The evaluation metrics are HD95 ($m$) and DSC in (%). DSC are reported for individual organs. $\uparrow\downarrow$ denotes the higher (lower) the better. - means missing data from the source. Bold and underline represent the best and the second best results. denotes that the experiment is conducted by us.
@@ IV Experiments
Zig-RiR (2D) [12] ($\times256$) 28.51 74.61 82.73 64.19 79.13 68.25 93.71 50.14 88.46 70.29
RWKV-UNet-T (ours) 11.89 81.87 88.10 68.42 88.88 82.61 95.15 60.52 90.23 81.09
RWKV-UNet-S (ours) 11.01 82.93 88.30 65.92 88.47 84.11 95.45 66.51 91.51 83.17
RWKV-UNet (ours) 08.85 84.29 88.71 70.39 89.58 84.36 95.70 66.58 93.54 85.49
TABLE IV: Comparison results on ACDC dataset. The evaluation metric is DSC in (%).
@@
Zig-RiR (2D) [12] ($\times256$) 87.42 86.41 81.87 93.98
RWKV-UNet-T (ours) 91.90 90.63 88.21 96.85
RWKV-UNet-S (ours) 91.19 89.49 87.87 96.22
RWKV-UNet (ours) 92.29 91.26 88.78 96.83
TABLE V: Comparison results on GOALS and FUGC 2025 dataset. The evaluation metric is DSC in (%). All experiments for baselines are conducted by us.
Methods GOALS FUGC 2025
U-Net [1] 90.99 51.96
TransUNet [28] 91.05 79.87
Rolling-UNet-M [22] 90.84 63.39
PVT-EMCAD-B2 [51] 90.91 57.39
PVTFormer [53] 90.98 78.41
UKAN [23] 90.59 72.19
VM-UNet [10] 91.10 72.94
H-vmunet [54] 84.18 65.99
Zig-RiR (2D) [12] 83.48 62.41
RWKV-UNet-T (ours) 91.34 78.67
RWKV-UNet-S (ours) 91.13 79.43
RWKV-UNet (ours) 91.75 81.17
@@
TABLE VI: Comparison results on BUSI, CVC-ClinicDB, CVC-ColonDB, Kvasir-SEG, ISIC 2017, GLAS and PSVFM datasets. The evaluation metric are DSC in (%). All experiments for baselines are conducted by us. Results with error bars are averaged over three runs with different splits.
@@
Methods BUSI CVC-ClinicDB CVC-ColonDB Kvasir-SEG ISIC 2017 GLAS PSVFM Params. FLOPs
[…]
Figure 4: Performance of different methods on the Synapse multi-organ segmentation dataset. The average DSC (%) is plotted against FLOPs (G). The size of each circle represents the model’s parameter count. RWKV-UNet achieves SOTA performance with balanced computation cost, while RWKV-UNet-S and RWKV-UNet-T also achieve remarkable results.
IV-A Experimental Setup
Datasets and Metrics
@@ IV-A Experimental Setup
Experiments on conducted on Synapse [56] for multi-organ Segmentation in CT images, ACDC [57] for cardiac segmentation in MRI images, GOALS [58] for OCT layer segmentation, BUSI [59] for breast tumor segmentation in ultrasound images, CVC-ClinicDB [60], CVC-ColonDB [61], Kvasir-SEG [62] for poly segmentation in endoscopy images, ISIC 2017 [63] for skin lesion segmentation in dermoscopic images, GLAS [64] for gland segmentation in microscopy images, PSVFM [65] for placental vessel segmentation in fetoscopy images and FUGC [66] for semi-supervised cervix segmentation in ultrasound images. The average Dice Similarity Coefficient (DSC) and average 95% Hausdorff Distance (HD95) are used as evaluation metrics.
  • Synapse uses 18 training and 12 validation CT scans, while ACDC uses 70 training, 10 validation, and 20 testing MRI scans, following settings in [28].
  • GOALS uses 100 training, 100 validation and 100 test images following the original setting in the challenge.
  • BUSI, CVC-ClinicDB, CVC-ColonDB, and Kvasir-SEG are split into training, validation, and testing sets with an 8:1:1 ratio, repeated with three different random seeds.
  • ISIC 2017 uses 2000 training, 150 validation, and 600 testing images as per the original challenge setup.
  • GlaS uses 85 training and 80 testing whole-slide images, with the training set further split into training and validation with a 9:1 ratio across three random seeds.
  • PSVFM uses a video-based split: Videos 1–3 for training, Video 4 for validation, and Videos 5–6 for testing, reflecting real-case variability in fetoscopic surgeries.
  • Building on [67] and [67], FUGC uses 50 labeled and 450 unlabeled images for training, 90 for validation, and 300 for testing.
Implementation Details
[…]
@@
where $\alpha$ and $\beta$ are 0.5 and 0.5 for Synapse, 0.2 and 0.8 for ACDC, 0.5 and 1 for GOALS and binary segmentation. Experiments for PVT-CASCADE [49] and PVT-EMCAD-B2 [51] use deep supervision strategy [69]. For semi-supervised tasks, the supervised loss on labeled data remains $\mathcal{L}_{\text{sup}}=0.5CE(\hat{y},y)+Dice(\hat{y},y)$. In addition, a consistency loss is applied on unlabeled data, computed as the mean squared error (MSE) between predictions under different augmentations after inverting the transformations. Learning rate for TransUNet and PVTFormer are $1e^{-4}$ on FUGC 2025 for better results.
IV-B Comparison with State-of-the-arts
Abdomen Multi-organ Segmentation
Table III shows that RWKV-UNet excels in abdomen organ segmentation on the Synapse dataset, achieving the highest average DSC of 84.29% , and surpassing all SOTA CNN-, transformer-, and Mamba-based methods. It also achieves 8.85 in HD95, which is significantly better than other models, demonstrating a strong ability to localize organ boundaries accurately. This can be attributed to the model’s ability to capture both long-range dependencies and local features. Smaller models, RWKV-UNet-T and RWKV-UNet-S, also achieve remarkable results. The DSCs with parameters and FLOPs of each model can be seen in Fig. 4 and visualization results can be seen in Fig. 5.
Cardiac Segmentation
Table IV shows that our model outperforms all SOTA models in the ACDC MRI cardiac segmentation task, achieving the best results in RV and the second-best in LV categories. The smaller models, RWKV-UNet-T and RWKV-UNet-S, also perform well.
OCT Layer Segmentation
Table V shows that on the small GOALS dataset, performance differences are minor and some carefully designed models even lag behind the classic U-Net, while our RWKV-UNet still achieves the highest average DSC, demonstrating its robustness and superiority. These results also reflect the model’s ability to handle small datasets.
Binary Segmentation
Table VI shows that the RWKV-UNet series shows remarkable performance in the tasks of breast lesion, poly, skin lesion, grand, and vessel segmentation, with much smaller parameters and computation load than TransUNet and FCBFormer.
Semi-supervised Cervix Segmentation
Table V also shows that the RWKV-UNet achieves SOTA performance on the FUGC 2025 dataset, further indicating strong data efficiency and generalization ability, as it can fully leverage limited labeled data and mine useful information from unlabeled data to maintain high segmentation accuracy.
@@ IV-B Comparison with State-of-the-arts
IV-C Ablation Study and Additional Analysis
Comparison with Different Encoders
We replace the encoder with pre-trained weights while keeping the rest of the network architecture unchanged (dimensions vary according to the encoder). Table VII shows that our pretrained encoder achieves the best DSC and HD95 on the Synapse dataset, with significantly lower FLOPs.
TABLE VII: Comparison results on Synapse dataset of different pretrained encoders
Encoder Average FLOPs
HD95 $\downarrow$ DSC $\uparrow$
ResNet 50 [45] 9.99 82.27 62.44
PVT-B3-v2 [70] 11.47 81.23 22.74
PVT-B5-v2 [70] 10.53 80.14 27.38
ConvNext-small [71] 12.79 76.34 42.47
ConvNext-base [71] 10.61 81.42 74.86
MaxViT-small [72] 10.57 83.71 18.54
MaxViT-base [72] 11.09 82.89 29.86
RWKV-UNet Enc-B (ours) 08.85 84.29 11.11
Effect of Pre-training for Encoder
Experimental results in Table VIII. indicate that pre-training is essential because it significantly improves feature extraction by enabling the encoder to capture intricate patterns and hierarchical information. And the inductive biases of RWKV are also not as strong as those of CNN, thus they also rely on pre-training similar to transformers.
TABLE VIII: Comparison results on Synapse dataset of whether to use pre-trained weights on ImageNet-1k or not.
@@ IV-C Ablation Study and Additional Analysis
w/o 30 36.23 71.66
w/o 100 25.01 78.14
w/ 30 08.85 84.29
TABLE IX: Comparison results of different decoder designs for RWKV-UNet on Synapse dataset.
Kernel Size Synapse ACDC FLOPs
HD95 $\downarrow$ DSC $\uparrow$ DSC $\uparrow$
3 10.35 84.83 92.07 10.97G
5 08.69 84.67 91.95 11.00G
7 09.97 83.73 91.61 11.05G
9 08.85 84.29 92.29 11.11G
11 11.25 82.67 92.04 11.19G
Attention in Shallow Stages
We evaluate the effect of retaining attention in the first two layers of RWKV-UNet-S by replacing them with spatial-mix attention and comparing against the original CNN-based shallow layers. All models are trained for 100 epochs from scratch, without pretraining. Table X shows negligible improvement from shallow attention on Synapse, while introducing higher computational cost, indicating that CNNs are sufficient for capturing low-level features in the shallow layers.
TABLE X: Comparison of results on the Synapse dataset: effect of attention in shallow stages of RWKV-UNet-S
Shallow Attention Average FLOPs
@@
w/ 29.43 76.88 6.91G
w/o (ours) 28.92 76.41 5.86G
Channel Mix in GLSP Blocks
In VRWKV [9], a Channel Mix module is used after the Spatial Mix. In contrast, our RWKV-UNet’s encoder designs GLSP blocks without Channel Mix, as the pointwise convolution in the output layer performs both channel mixing and feed-forward operations. We compare results on the Synapse dataset (without pre-training, 100 epochs) . Table XI demonstrates that Channel Mix is not essential for maintaining performance.
TABLE XI: Comparison of results on the Synapse dataset: effects of Channel Mix after Spatial Mix in GLSP blocks
@@
w/ 35.19 74.41 19.43G
w/o (ours) 25.01 78.14 11.11G
Skip Connections
Comparison results of different skip connection designs for RWKV-UNet on the Synapse dataset are shown in Table XII. Experiments analyze variations in the number of skip connections and the inclusion of the CCM block. We prioritize retaining shallow skip connections because shallow layers tend to experience greater information loss. The results show that, like other UNet’s variants, adequate skip connections are essential for RWKV-UNet’s performance. Furthermore, the results show that the CCM module can significantly improve the model segmentation results. With only two skips, the model can still achieve 84.28 DSC with CCM. Admittedly, the CCM module increases the computational load to some extent due to the aggregation of global information.
TABLE XII: Comparison results on different skip connections designs of RWKV-UNet on the Synapse dataset. Experiments are conducted on the number of skips and whether or not to use the CCM block.
@@
0 w 12.72 69.83 09.18 G
1 w/o 10.83 78.27 09.33 G
1 w/ 10.37 83.19 10.96 G
2 w/o 11.32 80.11 09.41 G
2 w/ 11.62 84.28 11.04 G
3 w/o 12.56 82.40 09.48 G
3 (ours) w/ (ours) 08.85 84.29 11.11 G
TABLE XIII: Comparison of different hidden rates in the ChannelMix of CCM on the Synapse dataset.
Hidden Rate Average FLOPs
@@
1 10.47 82.94 10.57 G
2 (ours) 8.85 84.29 11.11 G
3 9.89 84.18 11.62 G
4 9.24 84.06 12.14 G
Based on the results in Table XIII, the hidden rate of the Channel Mix layer in the CCM module materially affects performance: increasing the rate from 1 to 2 yields a notable improvement in performance, while further increases bring no gains at the cost of higher FLOPs.
Larger Resolution Inputs
The comparison results of different methods with $\times 512$ input in the Synapse dataset are shown in Table XIV. The average Dice scores of other methods are from [73]. Our results indicate that increasing the input resolution improves the performance of our model. Furthermore, compared to TransUNet, the computational overhead associated with the resolution increase is less severe, showcasing the efficiency of at higher resolutions.
TABLE XIV: Comparison of different methods with $\times512$ resolution input on the Synapse dataset. The evaluation metric is DSC in (%). The average DSCs of CNN-based methods are from [73].
@@
TransUNet [28] 84.36 148.29G
SAMed_h [76] 84.30 783.98G
RWKV-UNet (ours) 86.73 058.05G
V Conclusion
In this study, we introduce RWKV-UNet, a novel architecture that integrates the RWKV architecture with U-Net. By combining the strengths of convolutional networks for local feature extraction and RWKV’s ability to model global context, our model significantly improves medical image segmentation accuracy. The proposed enhancements, including the GLSP module and the CCM module, contribute to a more precise representation of features and information fusion across different scales. Experimental results on 11 datasets demonstrate that RWKV-UNet surpasses SOTA methods. Its variants (RWKV-UNet-S and RWKV-UNet-T) offer a practical balance between performance and computational efficiency. Our approach has a strong potential to advance medical image analysis, particularly in clinical settings where both accuracy and efficiency are paramount.
Limitations and Future Work. RWKV-UNet is a powerful 2D medical image segmentation model that effectively combines RWKV with convolutional operations; however, it is currently not applicable to 3D imaging. In the future, we plan to extend the model to 3D for handling volume data and to explore the potential of RWKV-based foundational models for medical imaging. We also aim to develop ultra-lightweight RWKV-based models tailored for point-of-care applications, preserving segmentation accuracy while enhancing adaptability and speed further.